Hello, I need help with this. E-mail me back as soon as possible, Thanks!
Fraction Action!
A man weighs five-sixth of his weight plus 30 pounds. What is his weight?

Found 2 solutions by ptaylor, josmiceli:
Answer by ptaylor(1894)

Let x=his weight
then (5/6)x+30=x multiply both sides by 6 and we get:
5x+180=6x subtract 5x from both sides
180=6x-5x and
x=180 lbs
ck(5/6)180+30=180
150+30=180
180=180
Hope this helps. Happy holidays----ptaylor

Answer by josmiceli(6785)

man's weight = w
w = (5/6)x + 30
w = x
x = (5/6)x + 30
(1/6)x = 30
x = 180 answer